### v1

# ColNet: Collaborative Optimization in Decentralized Federated Multi-task Learning Systems

Abstract
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has been explored to address client heterogeneity, with Federated Multi-Task Learning (FMTL) treating each client as a distinct task. However, most existing research focuses on data heterogeneity (e.g., addressing non-IID data) rather than task heterogeneity, where clients solve fundamentally different tasks. Additionally, much of the work relies on centralized settings with a server managing the federation, leaving the more challenging domain of decentralized FMTL largely unexplored. Thus, this work bridges this gap by proposing ColNet, a framework designed for heterogeneous tasks in decentralized federated environments. ColNet divides models into the backbone and task-specific layers, forming groups of similar clients, with group leaders performing conflict-averse cross-group aggregation. A pool of experiments with different federations demonstrated ColNet outperforms the compared aggregation schemes in decentralized settings with label and task heterogeneity scenarios.
1 Introduction
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has gained significant attention for its ability to tackle diverse but related tasks in a collaborative and privacy-preserving manner. FL trains models across clients by exchanging model updates rather than raw data, preserving data privacy Kairouz et al. (2021). Decentralized FL (DFL) eliminates the need for a central server, enhancing robustness by avoiding single points of failure Beltrán et al. (2023). MTL, traditionally performed on a single machine, leverages task interdependence to improve generalization, reduce overfitting, and address data sparsity Caruana (1997); Zhang and Yang (2021). Extending MTL to federated settings results in Federated Multi-Task Learning (FMTL) Smith et al. (2017), combining FL’s privacy preservation with MTL’s collaborative task-sharing benefits.
In both centralized and decentralized FMTL, existing research primarily addresses the challenge of non-independent and identically distributed (non-IID) data across clients, a core issue in FL Zhao et al. (2018). Non-IID data arises in various forms, including attribute skew (differences in feature distributions), label skew (differences in label distributions), and temporal skew (time-based changes in data distributions), each complicating model training Zhu et al. (2021). However, while much progress has been made in handling non-IID data, most studies focus on label heterogeneity rather than deeper task heterogeneity. Many approaches assume that all clients access the same set of class labels, even in non-IID settings. In reality, clients may encounter only subsets of labels relevant to their tasks Aoki et al. (2022), creating disparities in outputs and limiting collaborative learning. Addressing task heterogeneity, where clients solve fundamentally different problems, remains a significant challenge, especially in decentralized settings.
Aggregation approaches like FedAvg struggle in multi-task learning scenarios due to their reliance on model homogeneity Huang et al. (2022). Task heterogeneity introduces complexities, as clients may require distinct model architectures for tasks (e.g., classification versus point detection), leading to conflicting and dominant gradients during training Lu et al. (2024); Yu et al. (2024). While recent methods such as FedBone Chen et al. (2023) and FedHCA2 Lu et al. (2024) address these challenges, they focus on centralized settings (CFMTL), leaving decentralized FMTL (DFMTL) largely unexplored. Decentralization amplifies these challenges due to the absence of a coordinating server, demanding robust methods for managing tasks and model diversity.
This work introduces ColNet, a decentralized framework designed to train multi-task models in heterogeneous environments. ColNet groups clients by tasks, enabling intra-group federated averaging of backbones layers. Task group leaders then exchange averaged backbones across groups using a hyper-conflict-averse (HCA) aggregation scheme to reduce conflicting gradients and enhance cross-task collaboration.
Table 1: Overview of Research in Federated Multi-Task Learning
Federation Schema Techniques Research Approach used Heterogeneity Addressed
Convex MOCHA Smith et al. (2017) Convex optimization, dynamic client inclusion Label heterogeneity
Optimization OFMTL Li et al. (2019) Convex optimization, dynamic client inclusion Label heterogeneity
Regularization Ditto Li et al. (2021) Bi-level optimization, local regularization Label heterogeneity
PMTL Hu et al. (2021) Differential privacy, noise-injected global aggregation Label heterogeneity
Privatization FedPer Arivazhagan et al. (2019) Shared base layers, private personalization layers Label heterogeneity
Centralized FedRep Collins et al. (2021) Shared base layers, head freezing Label heterogeneity
Grouping CFL Sattler et al. (2020) Client grouping, cosine similarity Label heterogeneity
HeurFedAMP Huang et al. (2021) Dynamic grouping, cosine similarity Label heterogeneity
Knowledge EFDLS Xing et al. (2022) Knowledge distillation, local teacher-student models Label heterogeneity
Distillation FedICT Wu et al. (2023) Two-way knowledge distillation, reduced model updates Label heterogeneity
Hybrid FedBone Chen et al. (2023) Split learning, gradient projection, task adaptation Task heterogeneity
FedHCA2 Lu et al. (2024) Hyper conflict-averse aggregation, encoder-decoder architecture Task heterogeneity
Regularization DCLPMN Vanhaesebrouck et al. (2016) Graph-based regularization, neighbor collaboration Label heterogeneity
Decentralized Hybrid SpreadGNN He et al. (2022) Decentralized aggregation, molecular data focus Label heterogeneity
Hybrid ColNet (This work) Shared backbone, grouping, hyper conflict-averse aggregation Task heterogeneity
Experiments on CIFAR-10 and CelebA datasets demonstrate ColNet’s effectiveness. First, the framework’s performance is analyzed with varying aggregation round frequencies, showing that more frequent aggregations improve model performance. Second, different levels of layer privatization are examined, revealing that sharing most layers benefits class label heterogeneity setups, while increasing privatization slightly enhances performance in task heterogeneity scenarios. Finally, compared with other schemes, ColNet consistently outperforms alternatives in both class label and task heterogeneity scenarios, highlighting its capability to address diverse challenges in DFL.
2 Related Work
This section provides an overview of FMTL research, covering both centralized and decentralized approaches and comparing them based on the techniques employed and types of heterogeneity addressed, as presented in Table 1.
A. Centralized Approaches
FMTL was first introduced in 2017 with the MOCHA framework Smith et al. (2017), and since then, a variety of approaches have emerged to tackle different aspects of multi-task learning in centralized federated settings. Early work primarily focused on solving convex optimization problems, such as MOCHA and OFMTL Li et al. (2019), which were robust but unsuitable for modern non-convex deep learning methods. Subsequent research addressed this limitation, extending CFMTL to handle non-convex problems and introducing techniques like regularization, privatization, grouping, and knowledge distillation.
Ditto Li et al. (2021) introduced bi-level optimization techniques that allow each client to maintain a personalized model while regularizing it toward a global model. This approach ensures uniform performance across devices and enhances resistance to data and model poisoning attacks. Similarly, PMTL Hu et al. (2021) leverages relaxed differential privacy to learn personalized models while safeguarding client data. By adding noise during global aggregation, PMTL enables effective collaboration without compromising privacy. FedPer Arivazhagan et al. (2019) tackled statistical heterogeneity by dividing models into shared base layers and private personalization layers. Extensions like FedRep Collins et al. (2021) introduced mechanisms such as freezing the shared layers during early training phases to improve adaptability and achieve higher test accuracy.
Frameworks like CFL Sattler et al. (2020) introduced grouping strategies to cluster clients with similar data distributions, reducing the impact of conflicting gradients during aggregation. HeurFedAMP Huang et al. (2021) further refined this concept by using cosine similarity to dynamically determine grouping, allowing more granular adaptation to client data heterogeneity. EFDLS Xing et al. (2022) utilized knowledge distillation models to transfer knowledge efficiently within individual clients. This method enables collaborative learning by sharing only essential updates, reducing communication overhead. FedICT Wu et al. (2023) extended this idea by employing a two-way distillation process that customizes local models for specific tasks while maintaining a flexible and adaptive global model, making it suitable for diverse client environments.
Most existing studies primarily focus on class label heterogeneity, restricting their adaptability in dynamic environments. Frameworks like FedBone Chen et al. (2023) introduced split learning to separate general models on a central server from task-specific models on clients. Similarly, FedHCA2 Lu et al. (2024) proposed innovative aggregation techniques, such as HCA Aggregation, to address gradient conflicts, but it was designed for centralized environments.
B. Decentralized Approaches
DFMTL eliminates the need for a central server, allowing clients to communicate directly in a peer-to-peer manner. Despite the advantages, DFMTL introduces added challenges, such as managing gradient conflicts, asynchronous updates, and efficient client communication. Early research, such as DCLPMN Vanhaesebrouck et al. (2016), addressed these challenges by leveraging graph-based regularization to model task relationships, enabling collaboration between neighboring clients. SpreadGNN He et al. (2022) applied decentralized aggregation techniques tailored for tasks like molecular data analysis to improve scalability and address domain-specific requirements.
As presented in Table 1, while research on FMTL has made significant progress, most works addresses label heterogeneity and non-IID data rather than task heterogeneity, where clients solve fundamentally different tasks. Additionally, decentralized settings remain underexplored compared to centralized frameworks, despite their potential for improved privacy and scalability. Existing frameworks rarely address task heterogeneity in decentralized environments, leaving a clear gap for Heterogeneous DFMTL. This work addresses this gap by proposing a novel framework, ColNet, designed to handle both label and task heterogeneity in a decentralized setting.
3 The ColNet Solution
This work introduces ColNet, a collaborative optimization approach designed for multi-task learning in DFL. This section defines the DFMTL problem and outlines its optimization objectives. It then presents the proposed ColNet framework, describing each step in turn.
3.1 Problem Statement
In a DFMTL system with $N$ nodes, the learning objective is collaboratively learning task-specific models while preserving data privacy and operating under a fully decentralized communication topology.
Let $\mathcal{V}=\{1,2,\dots,N\}$ denote the set of nodes. Each node $i\in\mathcal{V}$ is associated with a task $\tau_{i}$ and holds a local dataset $\mathcal{D}_{i}$. The number of distinct tasks is $M\leq N$, accommodating the possibility that multiple nodes may share the same task or tasks may be heterogeneous across nodes.
The nodes form a connected, undirected graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{E}$ specifies the peer-to-peer (P2P) links. Communication is entirely decentralized, with no central server. Each node exchanges model updates or partial information solely with its neighbors $\mathcal{N}(i)$.
Each node $i$ maintains a parameter vector $\mathbf{w}_{i}\in\mathbb{R}^{d}$ that addresses its local task $\tau_{i}$. Heterogeneous local data render a single global model suboptimal, making personalized node-specific models more appropriate. Any common structure or correlation among tasks can be leveraged to improve overall performance.
Thus, the learning goal is framed as a joint optimization problem capturing both local performance and inter-task relationships:
$\min_{\{\mathbf{w}_{i}\}_{i=1}^{N}}\;\sum_{i=1}^{N}f_{i}\bigl{(}\mathbf{w}_{i};\mathcal{D}_{i}\bigr{)}\;+\;R\bigl{(}\{\mathbf{w}_{i}\}_{i=1}^{N}\bigr{)},$ (1)
where $f_{i}(\mathbf{w}_{i};\mathcal{D}_{i})$ is the local loss function at node $i$, and $R(\{\mathbf{w}_{i}\})$ is a regularization or coupling term designed to exploit shared structure among tasks.
3.2 ColNet Architecture
To address the challenges arising from multi-task learning in DFL, the ColNet framework leverages collaborative optimization to enable nodes to benefit from each other while handling diverse tasks. Figure 1 provides an overview of ColNet, which consists of two main phases: a pretraining phase and a DFL learning phase.
Figure 1: Overview of the ColNet Learning Process
Pretraining Phase
During the pretraining phase, nodes are clustered according to their data distributions and tasks. Nodes associated with the same task form a single task group. Within each task group, local models adopt the same architecture because they address an identical learning objective. In order to support multi-task learning, all nodes share a unified backbone network, whereas the top (task layers) differ depending on the particular task. Consequently, different task groups utilize a common backbone structure but maintain distinct task layers. Furthermore, each task group designates a task leader node, which serves as a communication bridge in subsequent cross-group information exchanges.
Learning Phase
Once task clustering is complete, the framework proceeds with the DFL learning process, which comprises four iterative steps, illustrated in Figure 1:
  1. 1.
    Local Training: Each node trains its local model on the corresponding dataset, updating both the backbone and task-layer parameters.
  2. 2.
    Intra-Task Aggregation: After a predefined number of local training epochs, an intra-task aggregation step is triggered. Nodes within the same task group exchange their updated backbone parameters and compute an averaged backbone across the group.
  3. 3.
    Cross-Task Aggregation: Each task leader node computes its backbone delta, the difference between the newly averaged backbone and the backbone used at the start of the local training phase, and shares this delta with the leaders of other task groups. Each task leader applies a simplified Hyper Conflict Averse Aggregation (HCA) to the received deltas, merging them into a new backbone.
  4. 4.
    Leader Aggregation and Distribution: The updated backbone is then propagated from leader nodes to all members of the corresponding task group for local model updates.
Upon completing these four steps, one training round of ColNet concludes, and the process repeats until a specified termination criterion is met. This iterative procedure allows each node to benefit from both intra-group collaborations with nodes sharing the same task and cross-group exchanges with nodes working on different tasks, thereby enhancing overall learning performance in a decentralized setting. The following subsections detail the steps in each phase.
3.3 Pretraining Phase
This section provides a detailed explanation of the Pretraining Phase in ColNet, covering both the task clustering process and the selection of leader nodes.
Task Clustering
A static grouping strategy is adopted for task clustering, where nodes assigned to the same task are placed in the same group. These nodes share an identical neural network architecture comprising both a backbone network and the task layers. Different task groups use the same backbone structure but employ different task layers to accommodate their respective tasks.
Leader Selection
A rotating-leader approach is adopted to distribute the leader’s computational overhead among group members. One member of each group is randomly selected as the leader node. After each cross-task group aggregation step, the current leader randomly selects a successor and notifies the corresponding client. The new leader then updates all other group leaders, who relay the information to their newly designated leaders.
3.4 Learning Phase
Compared to standard DFL methods, the main novelty of ColNet during training lies in introducing both intra-task and cross-task aggregation. The learning process in each client is presented in Algorithm 1.
Algorithm 1 ColNet Client Learning Process
1:Input: Intra-task peers $K$, leader set $L$, local epochs $epochs$, total rounds $rounds$, boolean $isLeader$
2:for $r=1$ to $rounds$ do
3:     for $e=1$ to $epochs$ do $\triangleright$ Local Training
4:         $\mathbf{w}\leftarrow$ Train locally
5:     end for
6:     for $k\in K$ do $\triangleright$ Intra-task Aggregation
7:         Send $\mathbf{w}^{B}$ to $k$
8:         Receive $\mathbf{w}^{B}_{k}$ from $k$
9:     end for
10:     $\mathbf{w}^{B}\leftarrow\text{Avg}\bigl{(}\{\mathbf{w}^{B}_{k}\}_{k\in K}\cup\{\mathbf{w}^{B}\}\bigr{)}$triangleright Eq. (2)
11:     if $isLeader$ then
12:         $\Delta\bar{\mathbf{w}}^{B}\leftarrow\mathbf{w}^{B}-\mathbf{w}^{B-1}$
13:         for $l\in L$ do $\triangleright$ Cross-task Aggregation
14:              Send $\Delta\bar{\mathbf{w}}^{B}$ to $l$
15:              Receive $\Delta\bar{\mathbf{w}}^{B}_{l}$ from $l$
16:         end for
17:         $\tilde{U}\leftarrow\text{HCA}\!\bigl{(}\Delta\bar{\mathbf{w}}^{B},\{\Delta\bar{\mathbf{w}}^{B}_{l}\}\bigr{)}$triangleright Eq. (4)
18:         $\mathbf{w}^{B}\leftarrow\mathbf{w}^{B}+\tilde{U}$
19:         for $k\in K$ do
20:              Send $\mathbf{w}^{B}$ to $k$
21:         end for
22:     else
23:         $\mathbf{w}^{B}\leftarrow\text{Wait for leader update}$
24:     end if
25:end for
Intra-Task Aggregation.
Within a task group, after each local training round, every node exchanges its updated backbone parameters with all other nodes in the same task group, then averages the collected backbones to update its own. Formally, if there are $K$ nodes in a group and $\mathbf{w}_{i}^{B}$ denotes the backbone parameters of node $i$, then the intra-task aggregation can be expressed as:
$\bar{\mathbf{w}}^{B}=\frac{1}{K}\sum_{i=1}^{K}\mathbf{w}^{B}_{i}.$ (2)
Cross-Task Aggregation.
Cross-task aggregation is more challenging due to potential conflicts among updates originating from distinct tasks. These conflicts arise when backbone parameters diverge significantly, leading to dominant or opposing gradients that degrade training. Inspired by FedHCA2 Lu et al. (2024), the cross-task aggregation in ColNet adopts a modified version of their gradient-alignment approach in decentralized learning environments.
ColNet uses a gradient term that maximizes the inner product between conflicting updates, effectively reducing gradient conflict. The core idea is to find an aggregated update $\tilde{U}$ (for the group leaders’ backbone) that remains close to the average update $\Delta\bar{\mathbf{w}}^{B}=\tfrac{1}{N}\sum_{i=1}^{N}\Delta\mathbf{w}i^{B}$ while minimizing inter-task conflicts:
$\max_{\tilde{U}}\min_{i}\langle\Delta\mathbf{w}_{i}^{B},\tilde{U}\rangle$ (3)
[…]
where $c\in[0,1)$ regulates how far $\tilde{U}$ can deviate from $\Delta\bar{\mathbf{w}}^{B}$.
By introducing a Lagrangian formulation and simplifying, the final update is shown to be:
$\tilde{U}\;=\;\Delta\bar{\mathbf{w}}^{B}\;+\;\frac{\sqrt{\phi}}{\|U_{w}\|}\,U_{w},$ (4)
where $U_{w}=\tfrac{1}{N}\sum_{i=1}^{N}w_{i}\Delta\mathbf{w}^{B}_{i}$ and $\phi=c^{2}\|\Delta\bar{\mathbf{w}}^{B}\|^{2}$. This formulation balances each leader’s update direction with the average update, mitigating conflicts while maintaining proximity to the mean. Upon computing $\tilde{U}$, each group leader updates its backbone and shares it with all group members, thereby propagating the conflict-reduced update throughout its task group.
4 Deployment and Evaluation
This section details the deployment and experimental evaluation of ColNet, covering the targeted multi-task learning scenarios, the chosen evaluation metrics, and the resulting performance outcomes.
4.1 Scenarios, Datasets, and Models
Two heterogeneity scenarios are considered for evaluation: label heterogeneity and task heterogeneity.
Label heterogeneity arises when clients train on different subsets of class labels, which requires splitting the model into a shared backbone and task-specific layers to protect privacy and maintain performance. Clients with the same labels form a single task group, sharing only the backbone parameters. The CIFAR-10 dataset Krizhevsky et al. (2009) was selected to evaluate label heterogeneity, where each image belongs to one of ten classes. Two task groups were formed: animals (bird, cat, deer, dog, frog, horse) and objects (airplane, automobile, ship, truck) classification. Each group contains three clients, and only the backbone parameters are shared within each group.
The dataset is split into 50,000 images for training and 10,000 for testing, further subdivided into animal or object sets. Each client’s training subset is then partitioned (80% train and 20% validation), and all clients use the same test subset in a task group. Data augmentation techniques, such as random cropping, flipping, and color jittering are applied to mitigate overfitting.
All clients train a ResNet-18 He et al. (2016) backbone and separate task layers to accommodate different output dimensions (four classes for objects, six for animals). A Cross-Entropy loss and SGD optimizer (learning rate 0.01, momentum 0.9, weight decay 1e-3) are employed.
Task heterogeneity is more complex, as clients may work on entirely different tasks (e.g., image classification and object detection), each requiring distinct final layers, loss functions, and optimizers. Dividing the model into a shared backbone and task-specific layer can facilitate aggregation in this scenario. The CelebA dataset Liu et al. (2015) is used to evaluate task heterogeneity. It contains over 200,000 face images with up to 40 binary attributes and five landmark annotations, making it suitable for multiple tasks, including face attribute recognition and landmark detection Woubie et al. (2024); Rasouli et al. (2020); Mortaheb et al. (2022).
Two distinct task groups are defined: one performs multi-label face attribute classification, and the other conducts face landmark detection. A subset of 40% of the dataset is split evenly between the groups (in total 80%), with each group’s clients receiving equal shares, and the remaining 20% subsets of data are used for testing. Clients then partition their data into 80% training and 20% validation while sharing a common test set per group. Random rotation and horizontal flips help combat overfitting.
Both tasks employ ResNet-18 backbone and distinct task layers:
  1. 1.
    Face attribute classification uses Binary Cross Entropy loss and an AdamW optimizer 5e-5 learning rate, 5e-3 weight decay.
  2. 2.
    Face landmark detection uses Mean Squared Error loss and an Adam optimizer 5e-4 learning rate.
The experiments report four evaluation metrics: loss, precision, recall, and F1-Score. Since face landmark detection is a regression task, only the loss metric is used to evaluate this subtask. To handle class imbalance and multi-class scenarios, micro averaging strategy is used, aggregating metrics across all classes within each task group. Metrics are kept separate for different groups because their labels (in the label heterogeneity scenario) or tasks (in the task heterogeneity scenario) are distinct. All models were developed and deployed in Python using the PyTorch framework, with Lightning as the trainer and GPUs for acceleration.
4.2 Federation Setup
For the federated setup, six nodes are split into two task groups, each fully connected internally, with inter-group communication handled by leader nodes. Fifteen training rounds are performed, each comprising two local epochs followed by intra-group and cross-group model aggregation. To investigate aggregation frequency, the experiment also compares 2, 3, and 5 local epochs per round.
Three baselines are implemented to assess the proposed ColNet framework:
  1. 1.
    No aggregation: Nodes train only on local data.
  2. 2.
    Intra-aggregation only: Nodes aggregate their backbone parameters within each group, with no cross-group exchange.
  3. 3.
    FedPre (intra- and cross-aggregation): Nodes use FedPre Arivazhagan et al. (2019) for both intra-group and cross-group backbone aggregation.
By contrast, ColNet combines FedPre for intra-group aggregation with HCA-based cross-group aggregation, enabling more effective handling of diverse tasks.
4.3 Impact of Aggregation Frequency
To maximize multi-task learning performance, it is important to determine the optimal aggregation frequency, i.e., how many local epochs to run per training round. Hence, the initial experiment compares ColNet’s performance with 2, 3, and 5 local epochs under two evaluation scenarios, thereby identifying the most suitable hyperparameter.
Table 2: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Local Epoch Setups
Scenario Subtask Epochs Loss Precision Recall F1-Score
Label Heterogeneity (CIFAR-10) Animal Classification 2 0.673 0.779 0.768 0.769
3 0.786 0.764 0.732 0.729
5 0.801 0.743 0.727 0.717
Object Classification 2 0.370 0.888 0.878 0.877
3 0.333 0.895 0.891 0.891
5 0.383 0.884 0.874 0.874
Task Heterogeneity (CelebA) Attribute Classification 2 0.231 0.741 0.545 0.605
3 0.233 0.735 0.542 0.598
5 0.237 0.733 0.522 0.579
Landmark Detection 2 5.299 - - -
3 5.224 - - -
5 5.762 - - -
Figure 2: Average Client Validation Loss in Each Epoch with Different Local Epoch Setups
Table 2 presents the results for four subtasks: animal classification, object classification, face attribute classification, and face landmark detection. Since face landmark detection is a regression task, only the loss metric is reported for that subtask. Besides, Figure 2 presents the epoch-wise validation loss for each subtask with different local training epoch settings.
In the label heterogeneity scenario (i.e., CIFAR-10), using two epochs per aggregation achieves the best performance for the animal group, while the object group slightly trails the three-epoch scheme. However, the difference is negligible, and both precision and recall remain closely aligned. From a training perspective (see Figure 2), the two-epoch setup converges more smoothly and exhibits lower loss. Despite the minor drop for the object group, the overall benefits of two-epoch intervals prevail.
Frequent aggregations benefit face attribute classification in the task heterogeneity scenario (i.e., CelebA). While the landmark detection task achieves its best results with three epochs, the two-epoch setting converges more quickly and smoothly. Consequently, two local training epochs per round are adopted in all remaining experiments.
4.4 Impact of the Privatization Degree
When deploying the backbone and task layers, a crucial hyperparameter concerns how to allocate these two parts of the network architecture. Although both scenarios in this work rely on ResNet-18 as the base model, determining which layers form the backbone and which form the task layers still requires careful consideration. Retaining more of the original, lower-level ResNet-18 layers in the backbone increases the amount of shared information among nodes, which may be detrimental for highly heterogeneous tasks. Conversely, expanding the upper residual layers into the task layer offers a higher degree of privatization but may weaken the effectiveness of aggregations.
Table 3: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Privatization Setups
Scenario Subtask Privatization Loss Precision Recall F1-Score
Label Heterogeneity (CIFAR-10) Animal Classification BL-0 0.706 0.786 0.759 0.758
BL-1 0.673 0.779 0.768 0.769
BL-2 0.810 0.739 0.720 0.716
Object Classification BL-0 0.313 0.900 0.895 0.895
BL-1 0.370 0.888 0.878 0.877
BL-2 0.371 0.881 0.876 0.875
Task Heterogeneity (CelebA) Attribute Classification BL-0 0.230 0.738 0.541 0.599
BL-1 0.231 0.741 0.545 0.605
BL-2 0.236 0.734 0.533 0.592
Landmark Detection BL-0 5.285 - - -
BL-1 5.299 - - -
BL-2 5.829 - - -
Figure 3: Average Client Validation Loss in Each Epoch with Different Privatization Setups
To evaluate how different levels of privatization impact multi-task learning, three configurations are compared: (i) no residual layers are privatized (BL-0), (ii) the last residual layer is part of the task layer (BL-1), and (iii) the last two residual layers are part of the task layer (BL-2). BL-0 represents the lowest degree of privatization, whereas BL-2 represents a high degree.
Table 3 compares the performance of ColNet client models under these three settings in both scenarios. Besides, Figure 3 presents the validation loss in each epoch during the training process. Overall, BL-2 shows the weakest performance, implying that retaining more residual layers in the backbone benefits DFMTL by providing sufficient neurons to exchange knowledge across different tasks. At the same time, a suitable number of task layers enables proper personalization under varied conditions. Drawing on the results from both scenarios, a moderate privatization strategy—treating the last residual block and the output layer of ResNet-18 as task layers (i.e., BL-1) proves optimal and is therefore adopted for subsequent experiments.
4.5 Comparison Among Aggregation Schemes
Having established an appropriate aggregation frequency and layer privatization strategy, the ColNet framework is now compared with other aggregation schemes. The main objective is to determine whether clients that leverage cross-task knowledge achieve better performance on their own tasks.
This experiment compares the performance of ColNet with (i) No aggregation (NO AGG.), (ii) Intra-aggregation (INTRA-AGG.), and (iii) FedPre with intra- and cross-aggregation (FEDPRE I&C), in label heterogeneity and task heterogeneity scenarios. Table 4 presents the test metrics of different aggregation schemes in these two scenarios after 15 federated rounds with a total of 30 epochs.
Table 4: Average Client Test Metrics for Different Aggregation Schemes
Scenario Subtask Aggregation Loss Precision Recall F1-Score
Label Heterogeneity (CIFAR-10) Animal Classification NO AGG. 0.938 0.699 0.675 0.669
INTRA-AGG. 0.953 0.723 0.680 0.679
FEDPRE(I&C) 0.853 0.730 0.700 0.689
ColNet 0.673 0.779 0.768 0.769
Object Classification NO AGG. 0.464 0.846 0.840 0.840
INTRA-AGG. 0.409 0.867 0.855 0.856
FEDPRE(I&C) 0.346 0.879 0.875 0.875
ColNet 0.370 0.888 0.878 0.877
Task Heterogeneity (CelebA) Attribute Classification NO AGG. 0.244 0.722 0.511 0.570
INTRA-AGG. 0.236 0.733 0.526 0.585
FEDPRE(I&C) 0.239 0.726 0.513 0.571
ColNet 0.231 0.741 0.545 0.605
Landmark Detection NO AGG. 6.054 - - -
INTRA-AGG. 5.392 - - -
FEDPRE(I&C) 5.325 - - -
ColNet 5.299 - - -
[NO AGG.: No aggregation; INTRA-AGG.: Intra-aggregation;
FEDPRE(I&C): FedPre with intra- and cross-aggregation]
Table 4 compares ColNet with three reference aggregation schemes under both the label heterogeneity and task heterogeneity scenarios. The results demonstrate that ColNet consistently outperforms the other methods across multiple datasets in multi-task settings.
In the label heterogeneity scenario, ColNet achieves the highest performance in both the animal and object classification subtasks. Notably, the F1-Score for animal classification exceeds the reference aggregation schemes by more than 0.08, signifying a substantial improvement. The findings further highlight that collaborative learning in a federated environment produces clearly superior results compared to no aggregation at all. Aggregating only within task groups increases precision by about 0.02, while employing an intra- and inter-group FedPer approach improves performance even more. Besides, ColNet’s cross-group HCA method effectively resolves conflicting gradients among heterogeneous tasks, allowing nodes to absorb knowledge from tasks beyond their own and enhancing local task performance.
A similar pattern is presented in the task heterogeneity scenario. Models without aggregation perform worst, whereas introducing inter-group aggregation yields a 0.01 gain in precision and a notable reduction in loss for the landmark detection subtask. With ColNet, both subtasks attain the best results: the F1-Score rises by 0.03, precision increases by 0.02, and the landmark detection loss drops by 0.8. These outcomes underscore ColNet’s clear advantages over other aggregation strategies in DFMTL.
Figure 4: Average Client Validation Loss in Each Epoch for Different Aggregation Schemes
Figure 4 presents the average client validation loss across epochs with using different aggregation schemes. Across all tasks, ColNet consistently achieves the lowest validation loss, highlighting its effectiveness in reducing errors during training. This advantage is particularly pronounced in animal classification and landmark detection, where ColNet’s smoother and more stable convergence clearly outperforms the other methods. The results underscore ColNet’s ability to manage both inter-task and intra-task heterogeneity more effectively than simpler aggregation schemes.
In contrast, FedPre and intra-aggregation show higher loss values and less stable convergence trends, while no aggregation performs the worst in all scenarios, reinforcing the critical importance of aggregation in federated multi-task learning. Notably, ColNet demonstrates a significant advantage by mitigating gradient conflicts and enabling knowledge sharing. These findings affirm ColNet’s superiority in DFMTL environments.
5 Conclusion and Future Work
This work introduces ColNet, a decentralized federated multi-task learning framework designed to handle both class label and task heterogeneity. ColNet organizes clients into task-specific groups, facilitating intra-group federated averaging of backbones layers. Leaders of each task group exchange the averaged backbones with leaders of other groups, employing an HCA aggregation scheme to mitigate conflicting gradients and promote effective cross-task collaboration.
Experiments across diverse datasets demonstrate its consistent superiority over other aggregation schemes. Besides, results indicate that higher aggregation frequency and a moderate degree of layer privatization could achieve performance gains.
Currently, ColNet is implemented on a single machine. Future research may investigate a fully decentralized deployment, including strategies for managing stragglers, faults, and malicious nodes, as well as dynamic clustering methods to replace predefined task groups. Additional comparisons with advanced algorithms under a truly decentralized setting could further refine ColNet’s capabilities.